0150 UT\xA0\xA0\xA0\xA0 Sunset. Unable to open due to high humidity.
0209 UT\xA0\xA0\xA0\xA0 Beginning calibration sequence (bias8k)
0230 UT\xA0\xA0\xA0\xA0 Beginning calibration sequence (bias1k)
0405 UT\xA0\xA0\xA0\xA0 Opening dome vents to equalize enclosure temperature. Cannot operate until T_mirror w/in 10C of outside temperature
0445 UT\xA0\xA0\xA0\xA0 Unable to open due to continuing high winds 1126 UT Twilight is beginning, so we are shutting down. No observations tonight.

-- David Atlee - 11 Apr 2013
Topic revision: r1 - 11 Apr 2013, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback